# On the Learn-to-Optimize Capabilities of Transformers in In-Context Sparse Recovery

### v2

@@ Abstract
An intriguing property of the Transformer is its ability to perform in-context learning (ICL), where the Transformer can solve different inference tasks without parameter updating based on the contextual information provided by the corresponding input-output demonstration pairs. It has been theoretically proved that ICL is enabled by the capability of Transformers to perform gradient-descent algorithms (Von Oswald et al., 2023a, ; Bai et al.,, 2024). This work takes a step further and shows that Transformers can perform learning-to-optimize (L2O) algorithms. Specifically, for the ICL sparse recovery (formulated as LASSO) tasks, we show that a $K$-layer Transformer can perform an L2O algorithm with a provable convergence rate linear in $K$. This provides a new perspective explaining the superior ICL capability of Transformers, even with only a few layers, which cannot be achieved by the standard gradient-descent algorithms. Moreover, unlike the conventional L2O algorithms that require the measurement matrix involved in training to match that in testing, the trained Transformer is able to solve sparse recovery problems generated with different measurement matrices. Besides, Transformers as an L2O algorithm can leverage structural information embedded in the training tasks to accelerate its convergence during ICL, and generalize across different lengths of demonstration pairs, where conventional L2O algorithms typically struggle or fail. Such theoretical findings are supported by our experimental results.
@@ 1 Introduction
Since its introduction in Vaswani et al., (2017), Transformers have become the backbone in various fields such as natural language processing (Radford,, 2018; Devlin,, 2018), computer vision (Dosovitskiy,, 2020) and reinforcement learning (Chen et al.,, 2021), significantly influencing subsequent research and applications. A notable capability of Transformers is their good performance for in-context learning (ICL) (Brown et al.,, 2020), i.e., without further parameter updating, Transformers can perform new inference tasks based on the contextual information embedded in example input-output pairs contained in the prompt. Such ICL capability facilitates state-of-the-art few-shot performances across a multitude of tasks, such as reasoning and language understanding tasks in natural language processing (Chowdhery et al.,, 2023), in-context dialog generation (Thoppilan et al.,, 2022) and in-context linear regression (Garg et al.,, 2022; Fu et al.,, 2023).
Given the significance of transformers’ ICL capabilities, extensive research efforts have been directed toward understanding the mechanisms behind their ICL performance. In this context, the ICL capability of a pre-trained Transformer is understood as the Transformer’s implicit implementation of learning algorithms during the forward pass. Von Oswald et al., 2023a , Dai et al., (2022), and Bai et al., (2024) suggest that these learning algorithms closely approximate gradient-descent-based optimizers, thus making the Transformer a universal solver for various ICL tasks. Specifically, these works demonstrate that Transformers can approximate gradient descent steps implicitly through some specific constructions of their parameters, enabling them to adapt to new data points during inference without explicit re-training.
However, it is well known that gradient-descent-based algorithms are not efficient in solving complicated optimization problems and thus may not be sufficient to explain the superior performance of Transformers on a plethora of ICL tasks. A recent work (Ahn et al.,, 2024) suggests that instead of gradient descent, Transformers actually perform pre-conditioned gradient descent during ICL. In other words, it learns a pre-conditioner during pre-training and then utilizes it during ICL to expedite the optimization process. Von Oswald et al., 2023b recently demonstrates that the forward pass of a trained transformer can implement meta-optimization algorithms, i.e., it can implicitly define internal objective functions and then optimize these objectives to generate predictions. Similarly, Zhang et al., (2023) show that a mesa-optimizer embedding the covariance matrix of input data can efficiently solve linear regression tasks. Such interpretation of the ICL mechanism shares the same essence as Learning-to-Optimize (L2O) algorithms, and motivates the following hypothesis:
@@
In this paper, we examine this hypothesis through the lens of in-context sparse recovery. Sparse recovery is a classical signal processing problem that is of significant practical interest across various domains, such as compressive sensing in medical imaging (Shen et al.,, 2017) and spectrum sensing (Elad,, 2010). Recent works show that Transformers are able to implement gradient descent-based algorithms with sublinear convergence rates for in-context sparse recovery (Bai et al.,, 2024; Chen et al., 2024b, ). However, empirical findings indicate that Transformers can solve in-context sparse recovery more efficiently than gradient descent-based approaches (Bai et al.,, 2024). Meanwhile, there exists a plethora of L2O algorithms that solve the classical sparse recovery problem efficiently with linear convergence guarantees (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019). Therefore, examining the L2O capabilities of Transformers in solving the in-context sparse recovery task becomes a promising direction and may serve as a perfect example to validate our hypothesis. Our main contributions are as follows.
@@
  • Second, we show that the Transformer as an L2O algorithm can actually outperform traditional L2O algorithms for sparse recovery in several aspects: 1) It does not require the measurement matrices involved in training to be the same, which is in stark contrast to traditional L2O algorithms (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019) for sparse recovery, and allows more flexibility to solve various in-context sparse recovery tasks. 2) It allows different numbers of measurements (i.e., prompt length) used for in-context sparse recovery, with guaranteed recovery performance as long as the number of measurements is sufficiently large. 3) It can extract structural properties of the underlying sparse vectors from the training data and utilize them to expedite its ICL convergence.
@@ 2 Related Works
Brown et al., (2020) first show that GPT-3, a Transformer-based LLM, can perform new tasks from input-output pairs without parameter updates, suggesting its ICL ability. This intriguing phenomenon of Transformers has attracted many attentions, leading to various interpretations and hypotheses about its underlying mechanism. For example, Han et al., (2023) empirically hypothesize that Transformers perform kernel regression with internal representations when facing in-context examples, and Fu et al., (2023) empirically show that Transformers learn to implement an algorithm similar to iterative Newton’s method for ICL tasks.
To better understand the ICL mechanism in large Transformers, existing works aim to demonstrate the Transformer’s capability for ICL by construction, e.g., showing that Transformers can perform gradient-based algorithms to solve ICL tasks by iteratively performing gradient descent layer by layer. In this category, Akyürek et al., (2022) show that by construction, Transformers can implement gradient descent-based algorithms for linear regression problems. Von Oswald et al., 2023a construct explicit weights for a Transformer, claiming it can perform gradient descent on linear and non-linear regression tasks. Bai et al., (2024) provide constructions such that Transformers can make selections between different gradient-based algorithms. Ahn et al., (2024) reason the ICL capability of Transformers to their ability to implement a pre-conditioned gradient descent for linear regression tasks, where the pre-condition matrix is learned from pre-training. Recently, Von Oswald et al., 2023b demonstrates that the forward pass of a trained transformer can implement meta-optimization algorithms, i.e., it can implicitly define internal objective functions and then optimize these objectives to generate predictions. Similarly, Zhang et al., (2023) show that a mesa-optimizer embedding the covariance matrix of input data can efficiently solve linear regression tasks. Our work belongs to this category, where we construct a Transformer structure that can implement an L2O algorithm to effectively solve the in-context sparse recovery problem.
@@
The L2O framework, as summarized in the review paper (Chen et al.,, 2022), is an optimization paradigm that develops an optimization method (i.e., a solver) by training across a set of similar problems (tasks) sampled from a task distribution. While the training process is often offline and time-consuming, the objective of L2O is to improve the optimization efficiency and accuracy when the method is deployed online and any new task sampled from the same distribution is encountered.
As a general optimization framework, L2O has demonstrated its advantage over classical static optimization frameworks in various optimization problems and applications. In this work, we focus on the L2O algorithms relevant to sparse recovery here and leave more discussions on general L2O in Appendix A. Sparse recovery, typically formulated as a least absolute shrinkage and selection operator (LASSO), has many important applications like magnetic resonance imaging (Meng et al.,, 2023) and stock market forecasting (Roy et al.,, 2015), thus motivates the design of efficient algorithms. E.g., the iterative soft-thresholding algorithm (ISTA) (Daubechies et al.,, 2004) is proposed to solve LASSO and improves over the standard gradient descent algorithm. Motivated by the ISTA structure, Gregor and LeCun, (2010) introduce the Learned ISTA (LISTA), a feedforward neural network that incorporates trainable matrices into ISTA updates. Chen et al., (2018) and Liu and Chen, (2019) further propose LISTA-Partial Weight Coupling (LISTA-CP) and Analytic LISTA (ALISTA) with fewer trainable parameters, making them easier to train. They also provide theoretical analyses demonstrating a linear convergence rate.
@@ 3 Preliminaries
For matrix $\mathbf{X}$, we use $[\mathbf{X}]_{p:q,r:s}$ to denote the submatrix that contains rows $p$ to $q$ and columns $r$ to $s$, and we use $[\mathbf{X}]_{:,i}$ and $[\mathbf{X}]_{j,:}$ to denote the $i$-th column and $j$-th row of $\mathbf{X}$ respectively. In some places, we also use $[\mathbf{X}]_{i}$ to denote its $i$-th column for convenience. We use $\|\mathbf{X}\|_{F}$ to denote its Frobenius norm. For vector $\mathbf{x}$, we use $\|\mathbf{x}\|_{1}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{\infty}$ to denote its $\ell_{1}$, $\ell_{2}$ and $\ell_{\infty}$ norms, respectively. We denote by $\mathbbm{1}_{d}$ and $\mathbf{0}_{d}$ the $d$-dimensional all-$1$ and all-00 column vectors, respectively. $\mathbbm{1}_{a\times b}$ and $\mathbf{0}_{a\times b}$ denote the all-$1$ and all-00 matrices of size $a\times b$, respectively.
@@ 3.1 Transformer Architecture
In this work, we consider the decoder-based Transformer architecture (Vaswani et al.,, 2017), where each attention layer is masked by a decoder-based attention mask and followed by a multi-layer perception (MLP) layer.
@@ 3.2 In-context Learning by Transformers
Previous studies about the mechanism of how Transformers performs in-context learning have attracted a lot of attention recently. To start with, it is believed that the ICL capability is due to the Transformer’s implicit implementation of learning algorithms in the forward pass. Von Oswald et al., 2023a , Dai et al., (2022), and Bai et al., (2024) suggest that these learning algorithms closely approximate gradient-descent-based optimizers, thus making the Transformer a universal solver for various ICL tasks. A recent work (Ahn et al.,, 2024) suggests that instead of gradient descent, Transformers actually perform pre-conditioned gradient descent for in-context least square linear regression. In general, these results corroborate the claim that the mechanism of the Transformer in-context learning is a L2O algorithm. We study this perspective and further provide the evidence supporting this claim by considering a more complicated in-context problem: in-context sparse recovery.
@@ 4.1 In-context Sparse Recovery
Sparse recovery is a fundamental problem in fields such as compressed sensing, signal denoising, and statistical model selection. The core concept of sparse recovery is that a high-dimensional sparse signal can be inferred from very few linear observations if certain conditions are satisfied. Specifically, it aims to identify an $S$-sparse vector $\bm{\beta}^{*}\in\mathbb{R}^{d}$ from its noisy linear observations $\mathbf{y}=\mathbf{X}\bm{\beta}^{*}+\bm{\epsilon}$, where $\mathbf{X}\in\mathbb{R}^{N\times d}$ is a measurement matrix, $\bm{\epsilon}\in\mathbb{R}^{N}$ is an isometric Gaussian noise vector with mean vector $\mathbf{0}_{N}$ and covariance matrix $\mathbf{I}_{N\times N}$. Typically, we assume $d\gg N$, which is the so-called under-determined case. One common assumption for the measurement matrix $\mathbf{X}$ (Pitaval et al.,, 2015; Ge et al.,, 2017; Zhu et al.,, 2021) is that each row of the matrix is independently sampled from an isometric sub-Gaussian distribution with zero mean and covariance matrix $\mathrm{diag}(\sigma_{1}^{2}\cdots,\sigma_{d}^{2})$, denoted as $P_{\mathbf{x}}$. which guarantees the critical restricted isometry property under mild conditions on the sparsity level (Candes and Tao,, 2007). In this work, we also assume $\bm{\beta}^{*}$ is randomly sampled from a distribution $P_{\bm{\beta}}$, which admits an $S$-sparse vector with random support.
@@ 4.2 Classical Algorithms
Gradient descent is known to struggle in solving the LASSO problem due to its inefficiency in effectively handling the sparsity constraint (Chen et al.,, 2018). This inefficiency has led to the development of more specialized algorithms that can better address the unique challenges posed by the LASSO formulation. A popular approach to solving the LASSO problem is the Iterative Shrinkage Thresholding Algorithm (ISTA). Starting with a fixed initial point $\bm{\beta}^{(1)}$, the update rule in the $k$-th iteration is given by
@@
Here, $\mathcal{S}_{\alpha/L}$ is the soft-thresholding function defined as $[\mathcal{S}_{\alpha/L}(\mathbf{x})]_{i}=\text{sign}([\mathbf{x}]_{i})\max\{0,|[\mathbf{x}]_{i}|-\alpha/L\}$, and $L$ is typically chosen as the largest eigenvalue of $\mathbf{X}^{\top}\mathbf{X}$ (Chen et al.,, 2018; Liu and Chen,, 2019).
Generally, for any ground-truth sparse vector $\bm{\beta}^{*}$ and any given $\mathbf{X}$, ISTA converges at a sublinear rate (Beck and Teboulle,, 2009). The sublinear convergence rate of ISTA is considered inefficient, which has led to the development of various LISTA-type L2O algorithms, such as LISTA (Gregor and LeCun,, 2010), LISTA-CP (Chen et al.,, 2018), and ALISTA (Liu and Chen,, 2019). These algorithms learn the weights in the matrices in ISTA rather than fixing them.
@@
where $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ are learnable parameters. Compared with ISTA with fixed parameters, LISTA-CP obtains $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ through pre-training. Specifically, with a fixed measurement matrix $\mathbf{X}$, it randomly samples $n$ $S$-sparse vectors $\{\bm{\beta}_{j}\}_{j=1}^{n}\sim P_{\beta}$ and generates $\{\mathbf{y}_{j}\}_{j=1}^{n}$, which is then utilized to optimize $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ by minimizing the total predicting loss for $\{\bm{\beta}_{j}\}_{j}$. Chen et al., (2018) show that, for the same measurement matrix $\mathbf{X}$, given any random instance $(\bm{\beta}^{*},\mathbf{y})$, a pre-trained LISTA-CP will converge to the ground-truth $\bm{\beta}^{*}$ linearly in $K$ under certain necessary conditions on ${\mathbf{X}}$.

### v3

@@ Abstract
An intriguing property of the Transformer is its ability to perform in-context learning (ICL), where the Transformer can solve different inference tasks without parameter updating based on the contextual information provided by the corresponding input-output demonstration pairs. It has been theoretically proved that ICL is enabled by the capability of Transformers to perform gradient-descent algorithms (vonoswald2023transformers; bai2023transformers). This work takes a step further and shows that Transformers can perform learning-to-optimize (L2O) algorithms. Specifically, for the ICL sparse recovery (formulated as LASSO) tasks, we show that a $K$-layer Transformer can perform an L2O algorithm with a provable convergence rate linear in $K$. This provides a new perspective explaining the superior ICL capability of Transformers, even with only a few layers, which cannot be achieved by the standard gradient-descent algorithms. Moreover, unlike the conventional L2O algorithms that require the measurement matrix involved in training to match that in testing, the trained Transformer is able to solve sparse recovery problems generated with different measurement matrices. Besides, Transformers as an L2O algorithm can leverage structural information embedded in the training tasks to accelerate its convergence during ICL, and generalize across different lengths of demonstration pairs, where conventional L2O algorithms typically struggle or fail. Such theoretical findings are supported by our experimental results.
@@ 1 Introduction
Since its introduction in vaswani2017attention, Transformers have become the backbone in various fields such as natural language processing (radford2018improving; devlin2019bert), computer vision (dosovitskiy2021image) and reinforcement learning (chen2021decision), significantly influencing subsequent research and applications. A notable capability of Transformers is their good performance for in-context learning (ICL) (brown2020language), i.e., without further parameter updating, Transformers can perform new inference tasks based on the contextual information embedded in example input-output pairs contained in the prompt. Such ICL capability facilitates state-of-the-art few-shot performances across a multitude of tasks, such as reasoning and language understanding tasks in natural language processing (chowdhery2022palm), in-context dialog generation (thoppilan2022lamda) and in-context linear regression (garg2023transformers; fu2023transformers).
Given the significance of transformers’ ICL capabilities, extensive research efforts have been directed toward understanding the mechanisms behind their ICL performance. In this context, the ICL capability of a pre-trained Transformer is understood as the Transformer’s implicit implementation of learning algorithms during the forward pass. vonoswald2023transformers, dai2023gpt, and bai2023transformers suggest that these learning algorithms closely approximate gradient-descent-based optimizers, thus making the Transformer a universal solver for various ICL tasks. Specifically, these works demonstrate that Transformers can approximate gradient descent steps implicitly through some specific constructions of their parameters, enabling them to adapt to new data points during inference without explicit re-training.
However, it is well known that gradient-descent-based algorithms are not efficient in solving complicated optimization problems and thus may not be sufficient to explain the superior performance of Transformers on a plethora of ICL tasks. A recent work (ahn2023transformers) suggests that instead of gradient descent, Transformers actually perform pre-conditioned gradient descent during ICL. In other words, it learns a pre-conditioner during pre-training and then utilizes it during ICL to expedite the optimization process. von2023uncovering recently demonstrates that the forward pass of a trained transformer can implement meta-optimization algorithms, i.e., it can implicitly define internal objective functions and then optimize these objectives to generate predictions. Similarly, zhang2023trained show that a mesa-optimizer embedding the covariance matrix of input data can efficiently solve linear regression tasks. Such interpretation of the ICL mechanism shares the same essence as Learning-to-Optimize (L2O) algorithms, and motivates the following hypothesis:
@@
In this paper, we examine this hypothesis through the lens of in-context sparse recovery. Sparse recovery is a classical signal processing problem that is of significant practical interest across various domains, such as compressive sensing in medical imaging (shen2017deep) and spectrum sensing (elad2010sparse). Recent works show that Transformers are able to implement gradient descent-based algorithms with sublinear convergence rates for in-context sparse recovery (bai2023transformers; chen2024transformers). However, empirical findings indicate that Transformers can solve in-context sparse recovery more efficiently than gradient descent-based approaches (bai2023transformers). Meanwhile, there exists a plethora of L2O algorithms that solve the classical sparse recovery problem efficiently with linear convergence guarantees (gregor2010learning; chen2018theoretical; liu2019alista). Therefore, examining the L2O capabilities of Transformers in solving the in-context sparse recovery task becomes a promising direction and may serve as a perfect example to validate our hypothesis. Our main contributions are as follows.
@@
  • Second, we show that the Transformer as an L2O algorithm can actually outperform traditional L2O algorithms for sparse recovery in several aspects: 1) It does not require the measurement matrices involved in training to be the same, which is in stark contrast to traditional L2O algorithms (gregor2010learning; chen2018theoretical; liu2019alista) for sparse recovery, and allows more flexibility to solve various in-context sparse recovery tasks. 2) It allows different numbers of measurements (i.e., prompt length) used for in-context sparse recovery, with guaranteed recovery performance as long as the number of measurements is sufficiently large. 3) It can extract structural properties of the underlying sparse vectors from the training data and utilize them to expedite its ICL convergence.
@@ 2 Related Works
brown2020language first show that GPT-3, a Transformer-based LLM, can perform new tasks from input-output pairs without parameter updates, suggesting its ICL ability. This intriguing phenomenon of Transformers has attracted many attentions, leading to various interpretations and hypotheses about its underlying mechanism. For example, han2023explaining empirically hypothesize that Transformers perform kernel regression with internal representations when facing in-context examples, and fu2023transformers empirically show that Transformers learn to implement an algorithm similar to iterative Newton’s method for ICL tasks.
To better understand the ICL mechanism in large Transformers, existing works aim to demonstrate the Transformer’s capability for ICL by construction, e.g., showing that Transformers can perform gradient-based algorithms to solve ICL tasks by iteratively performing gradient descent layer by layer. In this category, akyurek2022learning show that by construction, Transformers can implement gradient descent-based algorithms for linear regression problems. vonoswald2023transformers construct explicit weights for a Transformer, claiming it can perform gradient descent on linear and non-linear regression tasks. bai2023transformers provide constructions such that Transformers can make selections between different gradient-based algorithms. ahn2023transformers reason the ICL capability of Transformers to their ability to implement a pre-conditioned gradient descent for linear regression tasks, where the pre-condition matrix is learned from pre-training. Recently, von2023uncovering demonstrates that the forward pass of a trained transformer can implement meta-optimization algorithms, i.e., it can implicitly define internal objective functions and then optimize these objectives to generate predictions. Similarly, zhang2023trained show that a mesa-optimizer embedding the covariance matrix of input data can efficiently solve linear regression tasks. Our work belongs to this category, where we construct a Transformer structure that can implement an L2O algorithm to effectively solve the in-context sparse recovery problem.
@@
The L2O framework, as summarized in the review paper (chen2022learning), is an optimization paradigm that develops an optimization method (i.e., a solver) by training across a set of similar problems (tasks) sampled from a task distribution. While the training process is often offline and time-consuming, the objective of L2O is to improve the optimization efficiency and accuracy when the method is deployed online and any new task sampled from the same distribution is encountered.
As a general optimization framework, L2O has demonstrated its advantage over classical static optimization frameworks in various optimization problems and applications. In this work, we focus on the L2O algorithms relevant to sparse recovery here and leave more discussions on general L2O in Appendix A. Sparse recovery, typically formulated as a least absolute shrinkage and selection operator (LASSO), has many important applications like magnetic resonance imaging (meng2023artificial) and stock market forecasting (roy2015stock), thus motivates the design of efficient algorithms. E.g., the iterative soft-thresholding algorithm (ISTA) (daubechies2003iterative) is proposed to solve LASSO and improves over the standard gradient descent algorithm. Motivated by the ISTA structure, gregor2010learning introduce the Learned ISTA (LISTA), a feedforward neural network that incorporates trainable matrices into ISTA updates. chen2018theoretical and liu2019alista further propose LISTA-Partial Weight Coupling (LISTA-CP) and Analytic LISTA (ALISTA) with fewer trainable parameters, making them easier to train. They also provide theoretical analyses demonstrating a linear convergence rate.
@@ 3 Preliminaries
For matrix $\mathbf{X}$, we use $\mathbf{X}]_{p:q,r:s}$ to denote the submatrix that contains rows $p$ to $q$ and columns $r$ to $s$, and we use $\mathbf{X}]_{:,i}$ and $\mathbf{X}]_{j,:}$ to denote the $i$-th column and $j$-th row of $\mathbf{X}$ respectively. In some places, we also use $\mathbf{X}]_{i}$ to denote its $i$-th column for convenience. We use $\|\mathbf{X}\|_{F}$ to denote its Frobenius norm. For vector $\mathbf{x}$, we use $\|\mathbf{x}\|_{1}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{\infty}$ to denote its $\ell_{1}$, $\ell_{2}$ and $\ell_{\infty}$ norms, respectively. We denote by $\mathbbm{1}_{d}$ and $\mathbf{0}_{d}$ the $d$-dimensional all-11 and all-0 column vectors, respectively. $\mathbbm{1}_{a\times b}$ and $\mathbf{0}_{a\times b}$ denote the all-$1$ and all-0 matrices of size $a\times b$, respectively.
@@ 3.1 Transformer Architecture
In this work, we consider the decoder-based Transformer architecture (vaswani2017attention), where each attention layer is masked by a decoder-based attention mask and followed by a multi-layer perceptron (MLP) layer.
@@ 3.2 In-context Learning by Transformers
Previous studies about the mechanism of how Transformers performs in-context learning have attracted a lot of attention recently. To start with, it is believed that the ICL capability is due to the Transformer’s implicit implementation of learning algorithms in the forward pass. vonoswald2023transformers, dai2023gpt, and bai2023transformers suggest that these learning algorithms closely approximate gradient-descent-based optimizers, thus making the Transformer a universal solver for various ICL tasks. A recent work (ahn2023transformers) suggests that instead of gradient descent, Transformers actually perform pre-conditioned gradient descent for in-context least square linear regression. In general, these results corroborate the claim that the mechanism of the Transformer in-context learning is a L2O algorithm. We study this perspective and further provide the evidence supporting this claim by considering a more complicated in-context problem: in-context sparse recovery.
@@ 4.1 In-context Sparse Recovery
Sparse recovery is a fundamental problem in fields such as compressed sensing, signal denoising, and statistical model selection. The core concept of sparse recovery is that a high-dimensional sparse signal can be inferred from very few linear observations if certain conditions are satisfied. Specifically, it aims to identify an $S$-sparse vector $\bm{\beta}^{*}\in\mathbb{R}^{d}$ from its noisy linear observations $\mathbf{y}=\mathbf{X}\bm{\beta}^{*}+\bm{\epsilon}$, where $\mathbf{X}\in\mathbb{R}^{N\times d}$ is a measurement matrix, $\bm{\epsilon}\in\mathbb{R}^{N}$ is an isometric Gaussian noise vector with mean vector $\mathbf{0}_{N}$ and covariance matrix $\mathbf{I}_{N\times N}$. Typically, we assume $d\gg N$, which is the so-called under-determined case. One common assumption for the measurement matrix $\mathbf{X}$ (pitaval2015convergence; ge2017spurious; zhu2021global) is that each row of the matrix is independently sampled from an isometric sub-Gaussian distribution with zero mean and covariance matrix $\mathrm{diag}(\sigma_{1}^{2}\cdots,\sigma_{d}^{2})$, denoted as $P_{\mathbf{x}}$. which guarantees the critical restricted isometry property under mild conditions on the sparsity level (candes2007dantzig). In this work, we also assume $\bm{\beta}^{*}$ is randomly sampled from a distribution $P_{\bm{\beta}}$, which admits an $S$-sparse vector with random support.
@@ 4.2 Classical Algorithms
Gradient descent is known to struggle in solving the LASSO problem due to its inefficiency in effectively handling the sparsity constraint (chen2018theoretical). This inefficiency has led to the development of more specialized algorithms that can better address the unique challenges posed by the LASSO formulation. A popular approach to solving the LASSO problem is the Iterative Shrinkage Thresholding Algorithm (ISTA). Starting with a fixed initial point $\bm{\beta}^{(1)}$, the update rule in the $k$-th iteration is given by
@@
Here, $\mathcal{S}_{\alpha/L}$ is the soft-thresholding function defined as $\mathcal{S}_{\alpha/L}(\mathbf{x})]_{i}=\text{sign}([\mathbf{x}]_{i})\max\{0,|[\mathbf{x}]_{i}|-\alpha/L\}$, and $L$ is typically chosen as the largest eigenvalue of $\mathbf{X}^{\top}\mathbf{X}$ (chen2018theoretical; liu2019alista).
Generally, for any ground-truth sparse vector $\bm{\beta}^{*}$ and any given $\mathbf{X}$, ISTA converges at a sublinear rate (beck2009fast). The sublinear convergence rate of ISTA is considered inefficient, which has led to the development of various LISTA-type L2O algorithms, such as LISTA (gregor2010learning), LISTA-CP (chen2018theoretical), and ALISTA (liu2019alista). These algorithms learn the weights in the matrices in ISTA rather than fixing them.
@@
where $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ are learnable parameters. Compared with ISTA with fixed parameters, LISTA-CP obtains $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ through pre-training. Specifically, with a fixed measurement matrix $\mathbf{X}$, it randomly samples $n$ $S$-sparse vectors $\{\bm{\beta}_{j}\}_{j=1}^{n}\sim P_{\beta}$ and generates $\{\mathbf{y}_{j}\}_{j=1}^{n}$, which is then utilized to optimize $\{\theta^{(k)},\mathbf{D}^{(k)}\}$ by minimizing the total predicting loss for $\{\bm{\beta}_{j}\}_{j}$. chen2018theoretical show that, for the same measurement matrix $\mathbf{X}$, given any random instance $(\bm{\beta}^{*},\mathbf{y})$, a pre-trained LISTA-CP will converge to the ground-truth $\bm{\beta}^{*}$ linearly in $K$ under certain necessary conditions on $\mathbf{X}$}.